OpenStudy (anonymous):

Find the centripetal force needed by a 1250 kg car to make a turn of radius 40.0 m at a speed of 25.0 km/h.

OpenStudy (anonymous):

Centripetal force = mass*speed^2 /radius

OpenStudy (anonymous):

Okay that makes sense thank you

OpenStudy (anonymous):

i have the same problem but i plugged it in 1250*25^2/40 i got 19531.25 which is not one of my answers?

OpenStudy (anonymous):

@marr3000 Convert 25 km/h to m/s. So, 25km/h = 25*5/18 m/s and use this value for speed. :)

OpenStudy (anonymous):

wait so what would be the answer?

OpenStudy (anonymous):

@shailkumar ?

OpenStudy (anonymous):

Since centripetal force = mass*speed^2 /radius = \[\frac{1250 kg \times (25 \times5/18\ ms^{-1})^2 }{40m} \] = 217.01 N